# Sequential Multi-Object Grasping with One Dexterous Hand

### v1

@@ Abstract
Sequentially grasping multiple objects with multi-fingered hands is common in daily life, where humans can fully leverage the dexterity of their hands to enclose multiple objects. However, the diversity of object geometries and the complex contact interactions required for high-DOF hands to grasp one object while enclosing another make sequential multi-object grasping challenging for robots. In this paper, we propose SeqMultiGrasp, a system for sequentially grasping objects with a four-fingered Allegro Hand. We focus on sequentially grasping two objects, ensuring that the hand fully encloses one object before lifting it and then grasps the second object without dropping the first. Our system first synthesizes single-object grasp candidates, where each grasp is constrained to use only a subset of the hand’s links. These grasps are then validated in a physics simulator to ensure stability and feasibility. Next, we merge the validated single-object grasp poses to construct multi-object grasp configurations. For real-world deployment, we train a diffusion model conditioned on point clouds to propose grasp poses, followed by a heuristic-based execution strategy. We test our system using $8\times 8$ object combinations in simulation and $6\times 3$ object combinations in real. Our diffusion-based grasp model obtains an average success rate of 65.8% over 1600 simulation trials and 56.7% over 90 real-world trials, suggesting that it is a promising approach for sequential multi-object grasping with multi-fingered hands. Supplementary material is available on our project website: https://hesic73.github.io/SeqMultiGrasp.
@@ I Introduction
Recent efforts in dexterous grasping have leveraged dexterous hands for multi-object manipulation. For example, MultiGrasp [7] presents a framework for simultaneously grasping multiple objects using a Shadow Hand, and focuses on generating pre-grasp poses and executing lifting strategies. However, MultiGrasp targets scenarios where objects are positioned in close proximity, limiting its applicability and overlooking the potential of dexterous hands for independent control of multiple objects. Similarly, [10] explores multi-object grasping but assumes a fixed hand pose and requires human intervention to place objects within the hand. In contrast, we address the relatively unexplored domain of autonomous and sequential multi-object grasping, where a robot must secure one object, and then grasp the second object while maintaining control of the first.
Sequentially grasping multiple objects is particularly beneficial when objects are located too far apart for simultaneous manipulation. While one option might be to push objects together [11, 12] as a preprocessing step before simultaneous grasping, this might not always be feasible with high-friction surfaces and risks toppling certain objects.
@@
The contributions of this work are: (i) a novel framework, SeqMultiGrasp, for real-world sequential multi-object grasping; (ii) a new open-source dataset specifically designed to train sequential multi-object grasping methods; (iii) the first real-world robotic experiments demonstrating sequential multi-object grasping, where we conduct a total of 180 trials across 18 object pairs, with 90 trials replaying grasps from our dataset and 90 trials using grasp poses generated by our diffusion model.
@@ III Problem Statement and Assumptions
We study sequential multi-object grasping on a tabletop surface with $N_{o}$ objects, denoted as $\mathbf{O}=\{O_{j}\}_{j=1}^{N_{o}}$.
To represent each object $O_{j}$, we use a point cloud $P_{j}$ sampled from its surface $S(O_{j})$. To describe a single object, we can use $O$ and suppress the subscript when the distinction is not needed. Our objective is to generate a sequence of grasping actions $\mathcal{A}=\{a^{1},a^{2},\dots,a^{L}\}$ such that a robot sequentially grasps and lifts multiple objects while maintaining grasp stability for all objects in hand. We assume a robot arm with a single multi-fingered hand. In this work, we use a Franka Panda arm equipped with an Allegro Hand (see Fig. 1). We focus on when $N_{o}=2$, i.e., grasping two objects sequentially, and where the objects are sufficiently far apart that they cannot be grasped simultaneously with one hand. A trial is an instance of this sequential grasping problem, and is considered successful if both objects are lifted off the tabletop in sequence and can be stabilized for 2 seconds.
@@ IV-C Grasp Validation in Simulation
While the optimization procedure from Section IV-B creates grasp candidates, we still require validation to ensure that they lead to successful grasps. Thus, we execute grasps in a GPU-accelerated physics simulator, ManiSkill [39], to validate them while filtering out unsuccessful candidates.
@@
Rotation Robustness: From [3], we evaluate grasp stability under six different gravity directions. A grasp is considered valid if it maintains force closure across all tested conditions.
@@ IV-D Merging Multi-object Grasps
Figure 3: The objects we test in simulation (left) and real (right) for multi-object grasping. In both images, we use a dashed blue box to indicate the set of objects that the robot grasps first with the pinch grasp. Similarly, the dashed red box shows the set of objects that are grasped second with a side grasp. See Tables I and III for the name of these objects in simulation and real, respectively. The annotated numbers for simulation objects match those in Table I.
@@ IV-F Other Implementation Details
Instead, we adopt a heuristic-based approach inspired by prior work [4, 2], which avoids the complexities of RL. Specifically, we employ a simple squeeze-and-lift procedure. Given a hand configuration, we follow the strategy in Section IV-A to determine which fingers are used for each object. During grasp execution, we first use CuRobo [45] for motion planning to move the end effector to a collision-free pose set at a position offset from the grasp pose. Once in place, the end effector moves slowly toward the grasp pose without further collision checking.
@@ V Experiments
V-A Object Selection and Data Generation Evaluation
@@
To evaluate our dataset generation pipeline, we report the synthesis success rates for pinch-grasping objects and side-grasping objects (Table I). Due to the use of fewer hand links and our strict hybrid validation criteria, the success rates are lower compared to single-object grasping scenarios [3, 4, 2]. However, since the entire process is parallelized, this approach remains feasible despite the lower raw success rates. We store all resulting successful grasps in our data for subsequent experiments.
@@ V-B Simulation Experiment Setup
  • Dataset Replay (DR): We evaluate grasp poses sampled from our dataset. While each single-object grasp has already passed strict validation, failure may still occur after merging. This is mainly due to three factors: (1) Object placement shifts may lead to motion planning failures. (2) During execution, unintended interactions between the two objects or between the hand and objects may cause slippage or collisions. (3) A slight discrepancy exists between the validation and evaluation environments, as validation uses the PhysX GPU backend, while evaluation runs on the CPU.
@@
  • Ablation studies: We compare rotation representations for network prediction, including Euler angles, axis-angle, quaternion, 6D representation [49], and our chosen rotation matrix (orthogonalized via SVD). We test each of these representations over all $8\times 8$ object combinations using the LG experiment protocol.
Figure 4: A heat map where each grid cell shows the success rate (from 25 trials) of sequentially grasping two objects in simulation for dataset replay experiments. The first grasp is a pinch grasp (vertical axis) and the second grasp is a side grasp (horizontal axis). We refer to objects by an integer index (0 through 7). Table I lists the object names for each index.
@@ V-C Results from Simulation Experiments
In Fig. 4, we show a heatmap of the multi-object grasping success rate in simulation for our dataset replay experiments. Quantitatively, the results indicate that grasp poses from our dataset have potential for effective multi-object grasping. Most of the object-object grasping combinations result in a success rate above 80%, and four of them are at 100% (i.e., a 25 out of 25 success rate). The lowest observed success rate is 56%, occurring for object pairs involving the lemon, which is highly unstable on the tabletop and easily shifts upon contact with fingertips.
@@
Table II shows the result of our ablation study on the rotation representation choice for the hand pose. The results show that our rotation matrix representation results in the highest success rates, consistent with the findings of [4].
Euler Angle Axis Angle Quaternion 6D Ours
62.56% 60.25% 62.13% 64.25% 65.81%
TABLE II: Ablation study to compare rotation representations for hand poses. See Sections V-B and V-C for more details.
@@ V-D Real-World Experiment Setup
We conduct two types of real-world evaluations, based on Dataset Replay (DR) and Learned Grasp (LG) execution. These evaluations are similar to those in the simulation experiments (see Section V-B), except that for LG, the diffusion model is conditioned on real-world point clouds.
To reduce the sim-to-real gap, we follow the approach of [50] to obtain real-world object point clouds. To achieve this, we use RGB images of the entire scene from different camera views. We use Nerfstudio [51] to compute COLMAP reconstructions [52] and Stable Normal [53] to generate normal maps. Then, we employ 2D Gaussian Splatting [54] to obtain the point clouds. To determine the camera’s pose relative to the scene, we use AprilTag markers [55]. Given that we assume a canonical pose for the object, it is straightforward to crop out the object’s point cloud.
For both dataset replay and learned grasp execution, we conduct five trials per object pair. Since scene reconstruction is time-consuming, we perform it once per object combination. We use our trained diffusion model, conditioned on points sampled from the resulting real-world object mesh, to generate grasp poses.
@@ V-E Results from Real-World Experiments
We show an example successful trial in Fig. 1. See Fig. 5 for more example trials of SeqMultiGrasp. We report the results of our real-world experiments in Table III, which presents success rates for dataset replay (DR) and learned grasp execution (LG) across different object pairs. The average success rate for dataset replay is 64.4%, while the learned grasp execution achieves 56.7%.
@@
TABLE III: Real-world grasping results. “DR” refers to dataset replay, while “LG” refers to learned grasp execution using the diffusion model (see Section IV-E).
@@ ACKNOWLEDGMENTS
We thank Haozhe Lou, Zhiyuan Gao, Yiyang Ling, Chaoyi Pan, and Haoran Geng for their helpful technical advice and discussions. We are especially grateful to Yuyang Li for his invaluable help and support.

### v2

@@ Abstract
Sequentially grasping multiple objects with multi-fingered hands is common in daily life, where humans can fully leverage the dexterity of their hands to enclose multiple objects. However, the diversity of object geometries and the complex contact interactions required for high-DOF hands to grasp one object while enclosing another make sequential multi-object grasping challenging for robots. In this paper, we propose SeqMultiGrasp, a system for sequentially grasping objects with a four-fingered Allegro Hand. We focus on sequentially grasping two objects, ensuring that the hand fully encloses one object before lifting it and then grasps the second object without dropping the first. Our system first synthesizes single-object grasp candidates, where each grasp is constrained to use only a subset of the hand’s links. These grasps are then validated in a physics simulator to ensure stability and feasibility. Next, we merge the validated single-object grasp poses to construct multi-object grasp configurations. For real-world deployment, we train a diffusion model conditioned on point clouds to propose grasp poses, followed by a heuristic-based execution strategy. We test our system using $\times 8$ 8 object combinations in simulation and 6×36$\times$ 36 × 3 object combinations in real. Our diffusion-based grasp model obtains an average success rate of 65.8% over 1,600 simulation trials and 56.7% over 90 real-world trials, suggesting that it is a promising approach for sequential multi-object grasping with multi-fingered hands. Supplementary material is available on our project website: https://hesic73.github.io/SeqMultiGrasp.
@@ I Introduction
Recent efforts in dexterous grasping have leveraged dexterous hands for multi-object manipulation. For example, MultiGrasp [7] presents a framework for simultaneously grasping multiple objects using a Shadow Hand, and focuses on generating pre-grasp poses and executing lifting strategies. However, MultiGrasp targets scenarios where objects are positioned in close proximity, limiting its applicability and overlooking the potential of dexterous hands for independent control of multiple objects. Similarly, [10] explores multi-object grasping but assumes a fixed hand pose and requires human intervention to place objects within the hand. In contrast, we address the relatively unexplored domain of autonomous and sequential multi-object grasping, where a robot must secure one object, and then grasp the second object while maintaining control of the first. Sequentially grasping multiple objects is particularly beneficial when objects are located too far apart for simultaneous manipulation. While one option might be to push objects together [11, 12] as a preprocessing step before simultaneous grasping, this might not always be feasible with high-friction surfaces and risks toppling certain objects.
@@
The contributions of this work are: (i) a novel synthesis pipeline for generating physically-feasible multi-object grasp configurations; (ii) an integrated system, SeqMultiGrasp, for real-world sequential grasping of multiple objects with a dexterous hand; (iii) the first real-world robotic experiments demonstrating sequential multi-object grasping, where we conduct a total of 180 trials across 18 object pairs, with 90 trials directly executing synthesized grasp poses and 90 trials using grasp poses generated by our diffusion model.
@@ III Problem Statement and Assumptions
We study sequential multi-object grasping on a tabletop surface with $N_{o}$ objects, denoted as $\mathbf{O}=\{O_{j}\}_{j=1}^{N_{o}}$. To represent each object $O_{j}$, we use a point cloud $P_{j}$ sampled from its surface $S(O_{j})$. To describe a single object, we can use $O$ and suppress the subscript when the distinction is not needed. Our objective is to generate a sequence of $L$ grasping actions, $\mathcal{A}=\{a^{1},a^{2},\dots,a^{L}\}$, such that a robot sequentially grasps and lifts multiple objects while maintaining grasp stability for all objects in hand. We assume a robot arm with a single multi-fingered hand. In this work, we use a Franka Panda arm equipped with an Allegro Hand (see Fig. 1). We focus on when $N_{o}=2$, i.e., grasping two objects sequentially, and where the objects are sufficiently far apart that they cannot be grasped simultaneously with one hand. A trial is an instance of this sequential grasping problem, and is considered successful if both objects are lifted off the tabletop in sequence and can be stabilized for 2 seconds.
@@ IV-C Grasp Validation in Simulation
Rotation Robustness: From [3], we evaluate grasp stability under six axis-aligned gravity directions (±$\pm x,\pm y,\pm z$); a grasp is considered valid if the object remains in contact with the hand after 2.5 seconds of simulation across all tested conditions.
@@ IV-D Merging Multi-object Grasps
Figure 3: The objects we test in simulation (left) and real (right) for multi-object grasping. In both images, we use a dashed blue box to indicate the set of objects that the robot grasps first with the pinch grasp. Similarly, the dashed red box shows the set of objects that are grasped second with a side grasp. See Tables I and II for the name of these objects in simulation and real, respectively. The annotated numbers for simulation objects match those in Table I.
@@ IV-F Other Implementation Details
Instead, we adopt a heuristic-based approach inspired by prior work [4, 2], which avoids the complexities of RL. Specifically, we employ a simple squeeze-and-lift procedure. Although our execution procedure is not restricted to a specific finger-to-object assignment, we adopt the same strategy as in IV-A for consistency, as the diffusion model is biased toward the dataset’s distribution. During grasp execution, we first use CuRobo [45] for motion planning to move the end effector to a collision-free pose set at a position offset from the grasp pose. Once in place, the end effector moves slowly toward the grasp pose without further collision checking.
@@ V Experiments
V-A Object Selection and Grasp Synthesis Statistics
@@
To assess the efficiency of our single-object grasp pose synthesis method, we report the synthesis success rates for pinch-grasping objects and side-grasping objects (Table I). Due to the use of fewer hand links and our strict hybrid validation criteria, the success rates are lower compared to single-object grasping scenarios [3, 4, 2]. However, since the entire process is parallelized, this approach remains feasible despite the lower raw success rates. We store all resulting successful grasps in our data for subsequent experiments.
@@ V-B Simulation Experiment Setup
  • Synthesized Grasp (SG): We directly use grasp poses synthesized in the dataset for evaluation. Although each single-object grasp is strictly validated during dataset construction, failures may still occur. This is mainly due to: (1) object placement shifts leading to planning failures; (2) unintended interactions between the hand and objects or between the objects themselves; (3) a small simulation gap, as dataset validation is performed in batch on a GPU physics backend, while evaluation uses a CPU backend. This also serves as a baseline for the subsequent learned grasp evaluation.
@@
Figure 4: A heat map where each grid cell shows the success rate (from 25 trials) of sequentially grasping two objects in simulation for synthesized grasp experiments. The first grasp is a pinch grasp (vertical axis) and the second grasp is a side grasp (horizontal axis). We refer to objects by an integer index (0 through 7). Table I lists the object names for each index.
@@ V-C Results from Simulation Experiments
In Fig. 4, we show a heatmap of the multi-object grasping success rate in simulation for our synthesized grasp experiments. Quantitatively, the results indicate that grasp poses from our dataset have potential for effective multi-object grasping. Most of the object-object grasping combinations result in a success rate above 80%, and four of them are at 100% (i.e., a 25 out of 25 success rate). The lowest observed success rate is 56%, occurring for object pairs involving the lemon, which is highly unstable on the tabletop and easily shifts upon contact with fingertips.
@@ V-D Real-World Experiment Setup
We conduct two types of real-world evaluations, based on Synthesized Grasp (SG) and Learned Grasp (LG). These evaluations are similar to those in the simulation experiments (see Section V-B), except that for LG, the diffusion model is conditioned on real-world point clouds. To reduce the sim-to-real gap, we follow the approach of [49] to obtain real-world object point clouds. To achieve this, we use RGB images of the entire scene from different camera views. We use Nerfstudio [50] to compute COLMAP reconstructions [51] and Stable Normal [52] to generate normal maps. Then, we employ 2D Gaussian Splatting [53] to obtain the point clouds. To determine the camera’s pose relative to the scene, we use AprilTag markers [54]. Given that we assume a canonical pose for the object, it is straightforward to crop out the object’s point cloud.
For both synthesized and learned grasp evaluations, we conduct five trials per object pair. Since scene reconstruction is time-consuming, we perform it once per object combination. We use our trained diffusion model, conditioned on points sampled from the resulting real-world object mesh, to generate grasp poses.
@@ V-E Results from Real-World Experiments
We show an example successful trial in Fig. 1. See Fig. 5 for more example trials of SeqMultiGrasp. We report the results of our real-world experiments in Table II, which presents success rates for synthesized grasp (SG) and learned grasp (LG) across different object pairs. The average success rate for SG is 64.4%, while LG achieves 56.7%.
@@
TABLE II: Real-world grasping results. “SG” refers to synthesized grasp, while “LG” refers to learned grasp execution using the diffusion model (see Section IV-E).
@@ ACKNOWLEDGMENTS
We thank Haozhe Lou, Zhiyuan Gao, Yiyang Ling, Chaoyi Pan, and Haoran Geng for their helpful technical advice and discussions. We are especially grateful to Yuyang Li for his invaluable help and support. We sincerely thank the IROS 2025 reviewers for their constructive feedback.